OpenStudy (anonymous):

4. Describe how you would solve this equation for n: (−325.70) + (−27.90) + (−14.95) + 255 + n = 0

OpenStudy (idealist10):

-325.7-27.9-14.95+255+n=0 -113.55+n=0 n=0-(-113.55) n=0+113.55 n=113.55 Answer: 113.55